# Characterization of Gaussian Tensor Ensembles

### v1

@@ 3.2 Unitary case
Let us adapt this in the unitary case. Let $p$ be an even integer and $H$ a $p$-order hermitian tensor. The computation is slightly different from the one did for the case of local unitary invariance, where the legs are distinguished in $p$ different colors and cannot be matched with one of another color, see for instance the rich monograph of Gurau [7]. Let $R$ be an invariant polynomial of degree $n$. We still have $R(H)=\langle H^{\otimes p},C\rangle$, so for $U\in\mathrm{U}(N)$,
$R(H)=R(U^{T}\cdot H)=\langle U^{T}\cdot H^{\otimes n},C\rangle=\langle H^{\otimes n},U\cdot C\rangle=\langle H^{\otimes n},(U^{\otimes np/2}\otimes\bar{U}^{\otimes np/2})^{\sigma}\star_{np}C\rangle,$
[…]
$R(H)=\langle H^{\otimes n},\mathbb{E}_{U\in\mathrm{U}(N)}(U^{\otimes np/2}\otimes\bar{U}^{\otimes np/2})^{\sigma}\star_{np}C\rangle,$
@@
$(\mathbb{E}_{U\in\mathrm{U}(N)}(U^{\otimes np/2}\otimes\bar{U}^{\otimes np/2})^{\sigma})_{i_{1},\ldots,i_{2np}}=\sum_{\nu,\tau\in\mathrm{Match}(\{1,3,\ldots,np-1\},\{2,4,\ldots,np\})}\mathrm{Wg}_{\sigma,\tau}\prod_{t=1}^{np}\delta_{i_{t},i_{\nu(t)}}\delta_{i_{np+t},i_{np+\tau(t)}},$
where $\mathrm{Match}(\{1,3,\ldots,np-1\},\{2,4,\ldots,np\})$ is the set of involution with no fixed points from $\{1,\ldots,np\}$ into itself such that all the images of elements of $\{1,3,\ldots,np-1\}$ are in $\{2,4,\ldots,np\}$ and reciprocally. Hence,
$R(H)=\sum_{\tau\in\mathrm{Match}(\{1,3,\ldots,np-1\},\{2,4,\ldots,np\})}\alpha(\tau)\mathfrak{m}_{G(\tau)}(H),$
@@ 3.3 Symplectic case
$J=\mathbf{e_{2}}I^{\mathbb{H}}_{N}=\begin{pmatrix}0&-1&0&\ldots&0\\ 1&0&-1&\ldots&0\\ 0&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&-1\\ 0&\ldots&\ldots&1&0\end{pmatrix}=-J^{T}.$
@@
$R(H)=R(U^{T}\cdot H)=\langle H^{\otimes n},(U^{\otimes np/2}\otimes(-JUJ)^{\otimes np/2})^{\sigma}\star_{np}C\rangle,$
[…]
$R(H)=\langle H^{\otimes n},\mathbb{E}_{U\in\mathrm{Sp}(2N)}(U^{\otimes np/2}\otimes(-JUJ)^{\otimes np/2})^{\sigma}\star_{np}C\rangle,$
@@
(((( $\mathbb{E}_{U\in\mathrm{Sp}(2N)}(U^{\otimes np/2}\otimes(-JUJ)^{\otimes np/2})^{\sigma})_{i_{1},\ldots,i_{2np}}$
$=(-1)^{np/2}\sum_{k_{1},k^{\prime}_{2},\ldots,k_{np-1},k^{\prime}_{np}}\mathbb{E}_{U\in\mathrm{Sp}(2N)}\prod_{t=1}^{np/2}U_{i_{2t-1},i_{np+2t-1}}J_{i_{2t},k_{2t-1}}U_{k_{2t-1},k^{\prime}_{2t}}J_{k^{\prime}_{2t},i_{np+2t}}$
$=\sum_{k_{1},k^{\prime}_{2},\ldots,k_{np-1},k^{\prime}_{np}}\sum_{\nu,\tau\in\mathrm{Match}(\{1,3,\ldots,np-1\},\{2,4,\ldots,np\})}\mathrm{Wg}_{\nu,\tau}\prod_{t=1}^{np/2}J_{i_{2t-1},k_{\nu(2t)}}J_{i_{np+2t-1},k^{\prime}_{\tau(2t-1)}}J_{i_{2t},k_{2t-1}}J_{k^{\prime}_{2t},i_{np+2t}}$
$=\sum_{\nu,\tau\in\mathrm{Match}(\{1,3,\ldots,np-1\},\{2,4,\ldots,np\})}\mathrm{Wg}_{\nu,\tau}\prod_{t=1}^{np/2}\sum_{k}J_{i_{2t-1},k}J_{i_{\nu(2t-1)},k}\sum_{k^{\prime}}J_{i_{np+2t-1},k^{\prime}}J_{k^{\prime},i_{np+\tau(2t-1)}}$
$=\sum_{\nu,\tau\in\mathrm{Match}(\{1,3,\ldots,np-1\},\{2,4,\ldots,np\})}\mathrm{Wg}_{\nu,\tau}\prod_{t=1}^{np/2}\delta_{i_{2t-1},i_{\nu(2t-1)}}\delta_{i_{np+2t-1,i_{np+\tau(2t-1)}}}.$
$=\sum_{\nu,\tau\in\mathrm{Match}(\{1,3,\ldots,np-1\},\{2,4,\ldots,np\})}\mathrm{Wg}_{\nu,\tau}\prod_{t=1}^{np}\delta_{i_{t},i_{\nu(t)}}\delta_{i_{np+t,i_{np+\tau(t)}}}.$
@@
$R(H)=\sum_{\tau\in\mathrm{Match}(\{1,3,\ldots,np-1\},\{2,4,\ldots,np\})}\alpha(\tau)\mathfrak{m}_{G(\tau)}(H),$
@@ 3.4 The melon graph
The melon is the graph with two vertices linked by $p$ edges, in the hermitian case the $t$-th halfedge of the first vertex is linked with the $t+1$-th one of the second vertex and in the self-dual hermitian case the $2t$-th halfedge of a vertex is linked with the $12t-$ 1-th one of the other vertex. Then, the computation of the trace invariant is straightforward :
    in the real symmetric case,
    $\sum_{i_{1},\ldots,i_{p}}H_{i_{1},\ldots,i_{p}}H_{i_{1},\ldots,i_{p}}=\|H\|_{\mathrm{F}}^{2},$
@@
[…]
@@
    $\sum_{1\leq i_{1},\ldots,i_{p}\leq 2N}H_{i_{1},\ldots,i_{p}}H_{i_{2},i_{1},\ldots,i_{p},i_{p-1}}$ $=\sum_{i_{1},\ldots,i_{p}}(Q_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor})_{\iota_{1},\ldots,\iota_{p}}(Q_{\lfloor\frac{i_{2}}{2}\rfloor,\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor,\lfloor\frac{i_{p-1}}{2}\rfloor})_{\iota_{2},\iota_{1},\ldots,\iota_{p},\iota_{p-1}}$
    $=\sum_{i_{1},\ldots,i_{p}}\sum_{\epsilon\in\mathcal{E}}(Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor})^{2}(\otimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}})_{\iota_{1},\ldots,\iota_{p}}(\otimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}})_{\iota_{2},\iota_{1},\ldots,\iota_{p},\iota_{p-1}}$
    $\hskip 5.69046pt+\sum_{\epsilon\notin\mathcal{E}}(-1)^{p/2}(Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor})^{2}(\otimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}})_{\iota_{1},\ldots,\iota_{p}}(\otimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}})_{\iota_{2},\iota_{1},\ldots,\iota_{p},\iota_{p-1}}$
    $=\sum_{i_{1},\ldots,i_{p}}\sum_{\epsilon\in\mathcal{E}}(-1)^{n_{2}}(Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor})^{2}-\sum_{\epsilon\notin\mathcal{E}}(-1)^{n_{2}}(Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor})^{2},$
@@
    $\bar{H}_{i_{1},\ldots,i_{p}}$ $=\sum_{\epsilon:n_{1}\equiv n_{3}[2]}Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor}-\sum_{\epsilon:n_{1}\not\equiv n_{3}[2]}Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor}$
    $=\sum_{\epsilon:n_{1}\equiv n_{3}[2],n_{2}\equiv 0[2]}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}+\sum_{\epsilon:n_{1}\equiv n_{3}[2],n_{2}\equiv 1[2]}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}$
    $\hskip 28.45274pt-\sum_{\epsilon:n_{1}\not\equiv n_{3}[2],n_{2}\equiv 0[2]}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}-\sum_{\epsilon:n_{1}\not\equiv n_{3}[2],n_{2}\equiv 1[2]}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}$
    $=\sum_{\epsilon\in\mathcal{E}}(-1)^{n_{2}}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}-\sum_{\epsilon\notin\mathcal{E}}(-1)^{n_{2}}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\ldots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon},$
@@
    $\sum_{1\leq i_{1},\ldots,i_{p}\leq 2N}H_{i_{1},\ldots,i_{p}}H_{i_{2},i_{1},\ldots,i_{p},i_{p-1}}=\sum_{i_{1},\ldots,i_{p}}H_{i_{1},\ldots,i_{p}}\bar{H}_{i_{1},\ldots,i_{p}}=\|H\|_{\mathrm{F}}^{2}.$

### v2

@@ 3.2 Unitary case
Let us adapt this in the unitary case. Let $p$ be an even integer and $H$ a $p$-order hermitian tensor. The computation is slightly different from the one did for the case of local unitary invariance, where the legs are distinguished in $p$ different colors and cannot be matched with one of another color, see for instance the rich monograph of Gurau [7]. Let $R$ be an invariant polynomial of degree $n$. We still have $R(H)=\big\langle H^{\otimes p},C\big\rangle$, so for $U\in\mathrm{U}(N)$,
$R(H)$ $=R\bigl(U^{\mathsf{T}}\cdot H\bigr)=\big\langle U^{\mathsf{T}}\cdot H^{\otimes n},C\big\rangle$
$=\big\langle H^{\otimes n},U\cdot C\big\rangle=\big\langle H^{\otimes n},\bigl(U^{\otimes np/2}\otimes\bar{U}^{\otimes np/2}\bigr)^{\sigma}\star_{np}C\big\rangle,$
where we denote
  • $\sigma=(2\ \ np+1)(4\ \ np+3)\cdots(np\ \ 2np-1)$, and
  • $\bigl(U^{\otimes np/2}\otimes\bar{U}^{\otimes np/2}\bigr)^{\sigma}_{i_{1},\dots,i_{2np}}=\prod_{t=1}^{np/2}U_{i_{2t-1},i_{np+2t-1}}\bar{U}_{i_{2t},i_{np+2t}}$.
Taking the expectation over the Haar measure gives
$R(H)=\big\langle H^{\otimes n},\mathbb{E}_{U\in\mathrm{U}(N)}\bigl(U^{\otimes np/2}\otimes\bar{U}^{\otimes np/2}\bigr)^{\sigma}\star_{np}C\big\rangle,$
@@
$\bigl(\mathbb{E}_{U\in\mathrm{U}(N)}\bigl(U^{\otimes np/2}\otimes\bar{U}^{\otimes np/2}\bigr)^{\sigma}\bigr)_{i_{1},\dots,i_{2np}}$
$\qquad=\sum_{\nu,\tau\in\mathcal{M}(\{1,3,\dots,np-1\},\{2,4,\dots,np\})}\mathrm{Wg}_{\sigma,\tau}\prod_{t=1}^{np}\delta_{i_{t},i_{\nu(t)}}\delta_{i_{np+t},i_{np+\tau(t)}},$
where $\mathcal{M}(\{1,3,\dots,np-1\},\{2,4,\dots,np\})$ is the set of involution with no fixed points from $\{1,\dots,np\}$ into itself such that all the images of elements of $\{1,3,\dots,np-1\}$ are in $\{2,4,\dots,np\}$ and reciprocally ($\mathcal{M}$ is for “matching”). Hence,
$R(H)=\sum_{\tau\in\mathcal{M}(\{1,3,\dots,np-1\},\{2,4,\dots,np\})}\alpha(\tau)\mathfrak{m}_{G(\tau)}(H),$
@@ 3.3 Symplectic case
$J=\mathbf{e_{2}}I^{\mathbb{H}}_{N}=\begin{pmatrix}0&-1&0&\dots&0\\ 1&0&-1&\dots&0\\ 0&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&-1\\ 0&\dots&\dots&1&0\end{pmatrix}=-J^{\mathsf{T}}.$
@@
$R(H)=R\bigl(U^{\mathsf{T}}\cdot H\bigr)=\big\langle H^{\otimes n},\bigl(U^{\otimes np/2}\otimes(-JUJ)^{\otimes np/2}\bigr)^{\sigma}\star_{np}C\big\rangle,$
where we denote
  • $\sigma=(2\ \ np+1)(4\ \ np+3)\cdots(np\ \ 2np-1)$, and
  • $\bigl(U^{\otimes np/2}\otimes(-JUJ)^{\otimes np/2}\bigr)^{\sigma}_{i_{1},\dots,i_{2np}}=\prod_{t=1}^{np/2}U_{i_{2t-1},i_{np+2t-1}}(-JUJ)_{i_{2t},i_{np+2t}}$.
Again taking the expectation over the Haar measure gives
$R(H)=\big\langle H^{\otimes n},\mathbb{E}_{U\in\mathrm{Sp}(2N)}\bigl(U^{\otimes np/2}\otimes(-JUJ)^{\otimes np/2}\bigr)^{\sigma}\star_{np}C\big\rangle,$
@@
$\bigl(\mathbb{E}_{U\in\mathrm{Sp}(2N)}\bigl(U^{\otimes np/2}\otimes(-JUJ)^{\otimes np/2}\bigr)^{\sigma}\bigr)_{i_{1},\dots,i_{2np}}$
$\qquad=(-1)^{np/2}\sum_{k_{1},k^{\prime}_{2},\dots,k_{np-1},k^{\prime}_{np}}\mathbb{E}_{U\in\mathrm{Sp}(2N)}\prod_{t=1}^{np/2}U_{i_{2t-1},i_{np+2t-1}}J_{i_{2t},k_{2t-1}}U_{k_{2t-1},k^{\prime}_{2t}}J_{k^{\prime}_{2t},i_{np+2t}}$
$\qquad=\sum_{k_{i},k^{\prime}_{i}}\sum_{\nu,\tau\in\mathcal{M}(\{1,3,\dots,np-1\},\{2,4,\dots,np\})}\mathrm{Wg}_{\nu,\tau}$
$\phantom{\qquad=}{}\times\prod_{t=1}^{np/2}J_{i_{2t-1},k_{\nu(2t)}}J_{i_{np+2t-1},k^{\prime}_{\tau(2t-1)}}J_{i_{2t},k_{2t-1}}J_{k^{\prime}_{2t},i_{np+2t}}$
$\qquad=\sum_{\nu,\tau\in\mathcal{M}(\{1,3,\dots,np-1\},\{2,4,\dots,np\})}\mathrm{Wg}_{\nu,\tau}$
$\phantom{\qquad=}{}\times\prod_{t=1}^{np/2}\sum_{k}J_{i_{2t-1},k}J_{i_{\nu(2t-1)},k}\sum_{k^{\prime}}J_{i_{np+2t-1},k^{\prime}}J_{k^{\prime},i_{np+\tau(2t-1)}}$
$\qquad=\sum_{\nu,\tau\in\mathcal{M}(\{1,3,\dots,np-1\},\{2,4,\dots,np\})}\mathrm{Wg}_{\nu,\tau}\prod_{t=1}^{np/2}\delta_{i_{2t-1},i_{\nu(2t-1)}}\delta_{i_{np+2t-1,i_{np+\tau(2t-1)}}}.$
$\qquad=\sum_{\nu,\tau\in\mathcal{M}(\{1,3,\dots,np-1\},\{2,4,\dots,np\})}\mathrm{Wg}_{\nu,\tau}\prod_{t=1}^{np}\delta_{i_{t},i_{\nu(t)}}\delta_{i_{np+t,i_{np+\tau(t)}}}.$
@@
$R(H)=\sum_{\tau\in\mathcal{M}(\{1,3,\dots,np-1\},\{2,4,\dots,np\})}\alpha(\tau)\mathfrak{m}_{G(\tau)}(H),$
@@ 3.4 The melon graph
The melon is the graph with two vertices linked by $p$ edges, in the hermitian case the $t$-th halfedge of the first vertex is linked with the $(t+1)$-th one of the second vertex and in the self-dual hermitian case the $2t$-th halfedge of a vertex is linked with the (2t1)(2t-1)-th one of the other vertex. Then, the computation of the trace invariant is straightforward:
  • in the real symmetric case, $\sum_{i_{1},\dots,i_{p}}H_{i_{1},\dots,i_{p}}H_{i_{1},\dots,i_{p}}=\|H\|_{\mathrm{F}}^{2}$,
@@
    $\sum_{i_{1},\dots,i_{p}}H_{i_{1},\dots,i_{p}}H_{i_{2},\dots,i_{p},i_{1}}={}$ $\sum_{i_{1},\dots,i_{p}}\bigl(H^{(0)}_{i_{1},\dots,i_{p}}+iH^{(1)}_{i_{1},\dots,i_{p}}\bigr)$
    $\times\bigl(H^{(0)}_{i_{1},\dots,i_{p}}+i\underbrace{\epsilon((1\ \ 2\dots p))}_{=(-1)^{p-1}=-1}H^{(1)}_{i_{1},\dots,i_{p}}\bigr)=\|H\|_{\mathrm{F}}^{2},$
@@
    $\sum_{1\leq i_{1},\dots,i_{p}\leq 2N}$ $H_{i_{1},\dots,i_{p}}H_{i_{2},i_{1},\dots,i_{p},i_{p-1}}$
    $=\sum_{i_{1},\dots,i_{p}}\bigl(Q_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}\bigr)_{\iota_{1},\dots,\iota_{p}}\bigl(Q_{\lfloor\frac{i_{2}}{2}\rfloor,\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor,\lfloor\frac{i_{p-1}}{2}\rfloor}\bigr)_{\iota_{2},\iota_{1},\dots,\iota_{p},\iota_{p-1}}$
    $=\sum_{i_{1},\dots,i_{p}}\sum_{\epsilon\in\mathcal{E}}\bigl(Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}\bigr)^{2}\biggl(\bigotimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}}\biggr)_{\iota_{1},\dots,\iota_{p}}\biggl(\bigotimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}}\biggr)_{\iota_{2},\iota_{1},\dots,\iota_{p},\iota_{p-1}}$
    $\hskip 28.45274pt+\sum_{\epsilon\notin\mathcal{E}}(-1)^{p/2}\bigl(Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}\bigr)^{2}\biggl(\bigotimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}}\biggr)_{\iota_{1},\dots,\iota_{p}}\biggl(\bigotimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}}\biggr)_{\iota_{2},\iota_{1},\dots,\iota_{p},\iota_{p-1}}$
    $=\sum_{i_{1},\dots,i_{p}}\sum_{\epsilon\in\mathcal{E}}(-1)^{n_{2}}\bigl(Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}\bigr)^{2}-\sum_{\epsilon\notin\mathcal{E}}(-1)^{n_{2}}\bigl(Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}\bigr)^{2},$
@@
    $\bar{H}_{i_{1},\dots,i_{p}}$ $=\sum_{\epsilon\colon n_{1}\equiv n_{3}[2]}Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}-\sum_{\epsilon\colon n_{1}\not\equiv n_{3}[2]}Q^{\epsilon}_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}$
    $=\sum_{\epsilon\colon n_{1}\equiv n_{3}[2],\,n_{2}\equiv 0[2]}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}+\sum_{\epsilon\colon n_{1}\equiv n_{3}[2],\,n_{2}\equiv 1[2]}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}$
    $\hskip 28.45274pt-\sum_{\epsilon\colon n_{1}\not\equiv n_{3}[2],\,n_{2}\equiv 0[2]}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}-\sum_{\epsilon\colon n_{1}\not\equiv n_{3}[2],\,n_{2}\equiv 1[2]}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}$
    $=\sum_{\epsilon\in\mathcal{E}}(-1)^{n_{2}}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon}-\sum_{\epsilon\notin\mathcal{E}}(-1)^{n_{2}}Q_{\lfloor\frac{i_{1}}{2}\rfloor,\dots,\lfloor\frac{i_{p}}{2}\rfloor}^{\epsilon},$
@@
    $\sum_{1\leq i_{1},\dots,i_{p}\leq 2N}H_{i_{1},\dots,i_{p}}H_{i_{2},i_{1},\dots,i_{p},i_{p-1}}=\sum_{i_{1},\dots,i_{p}}H_{i_{1},\dots,i_{p}}\bar{H}_{i_{1},\dots,i_{p}}=\|H\|_{\mathrm{F}}^{2}.$
Acknowledgements